# Artificial Intelligence for Microbiology and Microbiome Research

### v1

@@ 3.2.2 Antibiotic resistance genes identification
Ground-breaking LLMs initially created for NLP have found success in predicting protein functions. These models, referred to as protein language models (PLMs), excel at generating intricate semantic representations that forge meaningful links between gene sequences and protein functions [62, 63, 64]. FunGeneTyper is a PLM-based deep learning framework designed for accurate and scalable prediction of protein-coding gene functions [142]. This framework includes two interconnected deep learning models: FunTrans and FunRep. While these models share a similar architecture, they are tailored for classifying functional genes at type and subtype levels, respectively. Both models utilize modular adapter-based architectures, incorporating a few additional parameters for efficient fine-tuning of extensive PLMs. Specifically, utilizing the ESM-1b model (a large-scale PLM built on a 33-layer transformer architecture [62]), adapters are inserted into each transformer layer, serving as individual modular units that introduce new weights tuned for specific tasks. FunGeneTyper has shown exceptional performance in classifying ARGs and virulence factor genes. More significantly, it is a flexible deep learning framework that can accurately classify general protein-coding gene functions and aid in discovering numerous valuable enzymes.
3.2.3 Plasmid identification
Plasmids are small, typically circular DNA molecules that are found in many microorganisms, e.g., Bactria, Archaea, and Eukaryota, which play an important role in microbial ecology and evolution through horizontal gene transfer, antibiotic resistance, and ecological interaction, etc. Identifying plasmid sequences from microbiome studies can provide a unique opportunity to study the mechanisms of plasmid persistence, transmission, and host specificity [143].
@@ 3.2.3 Plasmid identification
Many classical machine learning methods have been proposed for plasmid identification, e.g., cBar [144] based on sequential minimal optimization, PlasClass [145] using Logistic Regression, PlasmidVerify [146] using Naïve Bayesian classifier, PlasForest [147], Plasmer [148], Plasmidhunter [149], RFPlasmid [150] and SourceFinder [151] using Random Forest. Several deep learning methods have also been developed for plasmid identification. For example, PlasFlow employs MLP for the identification of bacterial plasmid sequences in environmental samples [152]. It can recover plasmid sequences from assembled metagenomes without any prior knowledge of the taxonomical or functional composition of samples with high accuracy. Deeplasmid is another deep learning method for distinguishing plasmids from bacterial chromosomes based on the DNA sequence [143]. It leverages both LSTM and fully connected layers to generate features, which are then concatenated and passed to another block of fully connected layers to generate the final output — the Deeplasmid score $y\in[0,1]$. The higher the score is for the sequence, the more likely it is to be a true plasmid. plASgraph2 is a new deep learning method for identifying plasmid contigs in fragmented genome assemblies built from short-read data [153]. The innovation of plASgraph2 lies in its use of GCN and the assembly graph to propagate information from neighboring nodes, resulting in more accurate classification. The GCN model consists of a set of graph convolutional layers designed to propagate information from neighboring contigs within the assembly graph. plASgraph2 generates two scores for each graph node: a plasmid score and a chromosomal score, which are used to assess whether a contig is likely derived from a plasmid, chromosome, or both.
Note that both plasmids and viruses are mobile genetic elements — a type of genetic material that can move around within a genome or be transferred from one species to another. Mobile genetic elements are often referred to as selfish genetic elements, because they have the ability to promote their own transmission at the expense of other genes in the genome. Mobile genetic elements are found in all organisms. The set of mobile genetic elements in an organism is called a mobilome, including viruses, plasmids, transposons, integrons, introns, etc. Recently, deep learning methods have been developed to simultaneously identify both viruses and plasmids, the two major components of the mobilome. For example, PPR-Meta is the first tool that can simultaneously identify phage and plasmid fragments from metagenomic assemblies efficiently and reliably [154]. PPR-Meta leveraged a novel architecture, Bi-path CNN, to improve the performance for short fragments. The Bi-path CNN leverages both base and codon information to enhance performance: the “base path” is effective for extracting sequence features of noncoding regions, while the “codon path” is useful for capturing features of coding regions. geNomad is a hybrid framework that combines the strengths of alignment-free and alignment-based models for concurrent identification and annotation of both plasmids and viruses in sequencing data [155]. To achieve that, geNomad processes user-provided nucleotide sequences via two distinct branches. In the sequence branch (“alignment-free”), the inputs are one-hot encoded and passed through an IGLOO neural network, which evaluates them by identifying non-local sequence motifs. In the marker branch (“alignment-based”), the proteins encoded by the input sequences are annotated with markers specific to chromosomes, plasmids, or viruses. Here, the key idea behind the IGLOO neural network is to leverage the relationships between “non-local patches” sliced from feature maps generated by successive convolutions to effectively represent long sequences, allowing it to handle both short and long sequences efficiently, unlike traditional RNNs which struggle with very long sequences [156].
@@ 3.2.4 Biosynthetic gene clusters prediction
Natural products are chemical compounds that serve as the foundation for numerous therapeutics in the pharmaceutical industry [157]. In microbes, these natural products are produced by clusters of colocalized genes known as biosynthetic gene clusters (BGCs) [158]. Advances in high-throughput sequencing have led to a surge in the availability of complete microbial isolate genomes and metagenomes, offering a great opportunity to discover a vast number of new BGCs. Deep learning models have been very useful in this genome mining effort [159, 160, 161, 162].
For example, DeepBGC and its extension employ (1) Pfam2vec (a word2vec-like word embedding model, which is a shallow neural network with a single hidden layer); (2) a Bidirectional LSTM (a classical RNN), which offers the advantage of capturing short- and long-term dependencies between adjacent and distant genes. e-DeepBGC still leverages those neural networks, but improves DeepBGC in the following aspects [159]. First, e-DeepBGC employs Pfam names, Pfam domain summary, Pfam domain clan information. This additional information is used to create new embedding of each Pfam domain by providing more biological information than that encoded by Pfam2vec which only uses the Pfam names. Second, a novel data augmentation step is introduced to overcome the limited number of BGCs with known functional classes.
BiGCARP is a self-supervised neural network masked language model [161]. It is based on the convolutional autoencoding representations of proteins (CARP), a masked language model of proteins. That’s why it is called Biosynthetic Gene CARP (or BiGCARP). The CARP is based on CNN, and has been shown to be competitive with transformer-based models for protein sequence pretraining [163]. SanntiS (Secondary metabolite gene cluster annotations using neural networks trained on InterPro signatures) is a new method for BGC prediction [164]. At the core of SanntiS is the detection model, a neural network with a one-dimensional convolutional layer, plus a bidirectional LSTM. This is quite similar to DeepBGC. The authors claimed that SanntiS outperforms DeepBGC, but it was not compared with BiGCAPR. Therefore, systematic benchmarking work is warranted.
@@ 3.3 Microbe-X Interactions
Recent advancements in microbiology and microbiome research have significantly deepened our understanding of the complex interactions between the microbes and the host, diseases, and drugs. In this section, we will discuss how deep learning-based methods have facilitated the inference of those complex interactions.
@@ 3.3.1 Microbe-host interactions
A disrupted gut microbiome has been linked to a wide variety of diseases, yet the mechanisms by which these microbes affect human health remain largely unclear. Protein-protein interactions (PPIs) are increasingly recognized as a key mechanism through which gut microbiota influence their human hosts [31, 176, 177, 178]. A vast and largely unexplored network of microbe-host PPIs may play a significant role in both the prevention and progression of various diseases. Future research is needed to further uncover these interactions and their potential therapeutic implications.
@@
Machine learning methods have been developed to study some of those mechanisms. For example, Morton et al. developed mmvec, a neural-network-based method to analyze microbe-metabolite interactions [187]. It takes microbial sequence counts and metabolite abundances from various samples as the input and outputs the estimated conditional probabilities of observing a metabolite given the presence of a specific microbe. This method is similar to a popular word embedding method in NLP, i.e., word2vec, which is a shallow neural network with a single hidden layer [188]. Note that in the original application of word2vec, the skip-gram technique (i.e., creating word embeddings that focus on predicting surrounding words based on a specific word or target word) was employed to account for the sequential nature of the text. For microbiome and metabolome data, there is no clear sequential nature. Therefore, in mmvec, the skip-gram was replaced by multinomial sampling, where a single microbe is randomly sampled from a microbiome sample at each gradient descent step. Morton et al. evaluated mmvec’s performance against traditional methods like Pearson’s, Spearman’s, SparCC, and SPIEC-EASI correlations, and found it demonstrated greater specificity and sensitivity, especially when applied to complicated datasets with vast amounts of microbiome and metabolomics information.
@@ 3.3.2 Microbe-disease associations
The exploration of microbe-disease associations (MDAs) is crucial for understanding various health conditions and tailoring effective treatments. Traditional studies directly correlate microbial features with disease outcomes, creating MDA databases such as HMDAD [189] and mBodyMap [190]. Advanced deep-learning methods have also been developed to infer new MDAs, including NinimHMDA [191], LGRSH [192], BPNNHMDA [193], and DMFMDA [194].
@@
Similar to NinimHMDA, LGRSH [192] and BPNNHMDA [193] were developed for the same predictive task but with different deep-learning architectures. LGRSH applies graph representation techniques to predict associations, using calculated similarities between microbes and diseases [192]. BPNNHMDA uses a back-propagation neural network to predict potential associations [193]. DMFMDA employs deep matrix factorization and Bayesian Personalized Ranking to predict associations [194]. Unfortunately, we haven’t seen any benchmark studies that systematically compare those deep learning methods in predicting microbiome-disease associations.
Very recently, thanks to the advancements in large language models, extraction of MDAs directly from biomedical literature has become much easier than before. For example, Karkera et al. demonstrated that pre-trained language models (specifically GPT-3, BioMedLM, and BioLinkBERT), when fine-tuned with domain and problem-specific data, can achieve state-of-the-art results for extracting MDAs from scientific publications [197]. The extracted MDAs will further expand the human MDA database. We expect that those deep learning methods will be more powerful with an expanded human MDA database.
@@ 3.4 Microbial Ecology
Deciphering inter-species interactions and assembly rules of microbial communities are fundamental but challenging questions in microbial ecology. Efforts based on population dynamics models have been made. However, parameterizing those dynamics models is very challenging  [210]. Deep learning approaches can overcome such challenges by learning the assembly rules implicitly without knowing the population dynamics. Especially, with the prominent progress in metagenomics and next-generation sequence technologies, collecting large-sample size data is feasible, providing sufficient diverse communities to train deep learning models.
3.4.1 Microbial interactions prediction
Microbes interact with each other and influence each other’s growth in various ways. The microbial interactions can be represented as a directed, signed, and weighted graph, i.e., the ecological network of the microbial community. Inferring the microbial interactions is important to understand the systems-level properties and dynamics of the microbial communities. Typically, this is achieved by analyzing high-quality longitudinal [211, 212, 213, 214, 215], or steady-state data [216], which is hard to obtain for large-scale microbial communities. Recently, the traditional random forest classifier was proposed to tackle this issue [217]. For each species, a trait is represented as a binary code in its trait vector. For each species pair within a community, a composite trait vector is created by concatenating the trait vectors of both species. This composite vector is then related to the observed responses of the interacting species. All interactions observed are utilized to train the classifier, which predicts the results of unobserved interactions. This approach has been evaluated in three case studies: a mapped interaction network of auxotrophic Escherichia coli strains, a soil microbial community, and a comprehensive in silico network illustrating metabolic interdependencies among 100 human gut bacteria. The results demonstrated that having partial knowledge of a microbial interaction network, combined with trait-level data of individual microbial species, can lead to accurate predictions of missing connections within the network, as well as propose potential mechanisms for these interactions. It would be very interesting to explore if deep learing methods can further improve the prediction of microbial interactions.
@@ 3.4.3 Keystone species identification
3.4.4 Colonization outcome prediction

### v2

@@ 3.2.2 Antibiotic resistance genes identification
Ground-breaking LLMs initially created for NLP have found success in predicting protein functions. These models, referred to as protein language models (PLMs), excel at generating intricate semantic representations that forge meaningful links between gene sequences and protein functions [60, 61, 62]. FunGeneTyper is a PLM-based deep learning framework designed for accurate and scalable prediction of protein-coding gene functions [275]. This framework includes two interconnected deep learning models: FunTrans and FunRep. While these models share a similar architecture, they are tailored for classifying functional genes at type and subtype levels, respectively. Both models utilize modular adapter-based architectures, incorporating a few additional parameters for efficient fine-tuning of extensive PLMs. Specifically, utilizing the ESM-1b model (a large-scale PLM built on a 33-layer transformer architecture [60]), adapters are inserted into each transformer layer, serving as individual modular units that introduce new weights tuned for specific tasks. FunGeneTyper has shown exceptional performance in classifying ARGs and virulence factor genes. More significantly, it is a flexible deep learning framework that can accurately classify general protein-coding gene functions and aid in discovering numerous valuable enzymes. ARG-BERT is another PLM-based model for predicting ARG resistance mechanisms [135]. Leveraging ProteinBERT [62] (a BERT-based PLM), ARG-BERT includes an input layer for ARG sequences, and an output layer for predicting six resistance mechanism labels: antibiotic efflux, antibiotic inactivation, antibiotic target alteration, antibiotic target protection, antibiotic target replacement, and others. ARG-BERT was fine-tuned using the HMD-ARG database [132] (composed of 17,282 high-quality sequences, coupled with labels of 15 antibiotic classes, 6 underlying resistance mechanisms, and their mobility), as well as a custom-built low-homology dataset. Subsequent attention analysis performed on the fine-tuned model (with HMD-ARG database) suggests that ARG-BERT takes into account biologically relevant features (e.g., antibiotic-binding sites and evolutionarily conserved regions) in making predictions.
3.2.3 Identification of mobile genetic elements
Mobile genetic elements are a type of genetic material that can move around within a genome or be transferred from one species to another. They are often referred to as selfish genetic elements, because they have the ability to promote their own transmission at the expense of other genes in the genome. They are found in all organisms. The set of mobile genetic elements in an organism is called a mobilome, including plasmids, viruses, transposons, integrons, introns, etc.
@@ 3.2.3 Identification of mobile genetic elements
Plasmids are small, typically circular DNA molecules that are found in many microorganisms. They play an important role in microbial ecology and evolution through horizontal gene transfer, antibiotic resistance, and ecological interaction, etc. Identifying plasmid sequences from microbiome studies can provide a unique opportunity to study the mechanisms of plasmid persistence, transmission, and host specificity [137]. Many classical machine learning methods have been proposed for plasmid identification, e.g., cBar [276] based on sequential minimal optimization, PlasClass [277] using Logistic Regression, PlasmidVerify [278] using Naïve Bayesian classifier, PlasForest [279], Plasmer [280], Plasmidhunter [281], RFPlasmid [282] and SourceFinder [283] using Random Forest. Several deep learning methods have also been developed for plasmid identification. For example, PlasFlow employs MLP for the identification of bacterial plasmid sequences in environmental samples [136]. It can recover plasmid sequences from assembled metagenomes without any prior knowledge of the taxonomical or functional composition of samples with high accuracy. Deeplasmid is another deep learning method for distinguishing plasmids from bacterial chromosomes based on the DNA sequence [137]. It leverages both LSTM and fully connected layers to generate features, which are then concatenated and passed to another block of fully connected layers to generate the final output–the Deeplasmid score. The higher the score is for the sequence, the more likely it is to be a true plasmid. plASgraph2 is a new deep learning method for identifying plasmid contigs in fragmented genome assemblies built from short-read data [138]. The innovation of plASgraph2 lies in its use of GCN and the assembly graph to propagate information from neighboring nodes, resulting in more accurate classification. The GCN model consists of a set of graph convolutional layers designed to propagate information from neighboring contigs within the assembly graph. plASgraph2 generates two scores for each graph node: a plasmid score and a chromosomal score, which are used to assess whether a contig is likely derived from a plasmid, chromosome, or both.
Recently, deep learning methods have been developed to simultaneously identify both viruses and plasmids, the two major components of the mobilome. For example, PPR-Meta is the first tool that can simultaneously identify phage and plasmid fragments from metagenomic assemblies efficiently and reliably [139]. PPR-Meta leveraged a novel architecture, Bi-path CNN, to improve the performance for short fragments. The Bi-path CNN leverages both base and codon information to enhance performance: the “base path” is effective for extracting sequence features of noncoding regions, while the “codon path” is useful for capturing features of coding regions.
@@ 3.2.4 Biosynthetic gene clusters prediction
Natural products are chemical compounds that serve as the foundation for numerous therapeutics in the pharmaceutical industry [284]. In microbes, microbial secondary metabolites (an important source of natural products) are produced by clusters of colocalized genes known as biosynthetic gene clusters (BGCs) [285]. Advances in high-throughput sequencing have led to a surge in the availability of complete microbial isolate genomes and metagenomes, offering a great opportunity to discover a vast number of new BGCs. BGC identification primarily employs rule-based or machine learning approaches. Rule-based methods (e.g., antiSMASH [286] and PRISM [287]) excel at detecting known BGC categories (e.g., alkaloids, nonribosomal peptides, polyketides, ribosomally synthesized and post-translationally modified peptides (RiPP), saccharides, terpenes) but struggle with novel BGCs. Many machine learning methods [288, 289, 290, 140, 141, 142, 143] have been developed to overcome the limitations of rule-based methods. In particular, deep learning techniques have been very useful in this endeavor.
For example, DeepBGC and its extension employ (1) Pfam2vec (a word2vec-like word embedding model, which is a shallow neural network with a single hidden layer); (2) a Bidirectional LSTM (a classical RNN), which offers the advantage of capturing short- and long-term dependencies between adjacent and distant genes [140]. e-DeepBGC still leverages those neural networks, but improves DeepBGC in the following aspects [141]. First, e-DeepBGC employs Pfam names, Pfam domain summary, Pfam domain clan information. This additional information is used to create new embedding of each Pfam domain by providing more biological information than that encoded by Pfam2vec which only uses the Pfam names. Second, a novel data augmentation step is introduced to overcome the limited number of BGCs with known functional classes.
BiGCARP is a self-supervised neural network masked language model [142]. It is based on the convolutional autoencoding representations of proteins (CARP), a masked language model of proteins. That’s why it is called Biosynthetic Gene CARP (or BiGCARP). The CARP is based on CNN, and has been shown to be competitive with transformer-based models for protein sequence pretraining [291].
BGC-Prophet is a transformer-based language model for BGC prediction and classification [143]. It employs ESM2 for sequence-specific gene embeddings and a transformer encoder with classifiers for BGC detection and product classification. This approach enhances accuracy and efficiency in identifying both known and novel BGCs by capturing location-dependent gene relationships. In particular, BGC-Prophet is several orders of magnitude faster than existing methods, e.g., DeepBGC, enabling pan-phylogenetic screening and whole-metagenome screening of BGCs.
@@ 3.3 Microbe-X Interactions
Recent advancements in microbiology and microbiome research have significantly deepened our understanding of the complex microbial interactions, as well as interactions between the microbes and the host, diseases, and drugs. In this section, we will discuss how deep learning-based methods have facilitated the inference of those complex interactions.
3.3.1 Microbe-microbe interactions
Microbes interact with each other and influence each other’s growth in various ways. Inferring the microbial interactions is important to understand the systems-level properties of the microbial communities. Typically, this is achieved by analyzing high-quality longitudinal [301, 302, 303, 304, 305], or steady-state data [306], which is hard to obtain for large-scale microbial communities. Recently, the traditional random forest classifier was proposed to tackle this issue [146]. For each species, a trait is represented as a binary code in its trait vector. For each species pair within a community, a composite trait vector is created by concatenating the trait vectors of both species. This composite vector is then related to the observed responses of the interacting species. All interactions observed are utilized to train the classifier, which predicts the results of unobserved interactions. This approach has been evaluated in three case studies: a mapped interaction network of auxotrophic Escherichia coli strains, a soil microbial community, and a comprehensive in silico network illustrating metabolic interdependencies among 100 human gut bacteria. The results demonstrated that having partial knowledge of a microbial interaction network, combined with trait-level data of individual microbial species, can lead to accurate predictions of missing connections within the network, as well as propose potential mechanisms for these interactions. It would be very interesting to explore if deep learing methods can further improve the prediction of microbial interactions.
@@ 3.3.2 Microbe-host interactions
A disrupted gut microbiome has been linked to a wide variety of diseases, yet the mechanisms by which these microbes affect human health remain largely unclear. Protein-protein interactions (PPIs) are increasingly recognized as a key mechanism through which gut microbiota influence their human hosts [31, 307, 308, 309]. A vast and largely unexplored network of microbe-host PPIs may play a significant role in both the prevention and progression of various diseases. Future research is needed to uncover these interactions and their potential therapeutic implications.
@@ 3.3.3 Microbe-disease associations
The exploration of microbe-disease associations (MDAs) is crucial for understanding various health conditions and tailoring effective treatments. Traditional studies directly correlate microbial features with disease outcomes, creating MDA databases such as HMDAD [313] and mBodyMap [314]. Advanced deep-learning methods have also been developed to infer new MDAs. Those methods include NinimHMDA [152], LGRSH [153], BPNNHMDA [154],
DMFMDA [315], GATMDA [316], MVGCNMDA [317], KGNMDA [318], GPUDMDA [319] and GCATCMDA [320].
@@
Similar to NinimHMDA, LGRSH [153] and BPNNHMDA [154] were developed for the same predictive task but with different deep-learning architectures. LGRSH applies graph representation techniques to predict associations, using calculated similarities between microbes and diseases [153]. BPNNHMDA uses a back-propagation neural network to predict potential associations [154]. DMFMDA employs deep matrix factorization and Bayesian Personalized Ranking to predict associations [315].
Thanks to the advancements in large language models, extraction of MDAs directly from biomedical literature has become much easier than before. For example, Karkera et al. demonstrated that pre-trained language models (specifically GPT-3, BioMedLM, and BioLinkBERT), when fine-tuned with domain and problem-specific data, can achieve state-of-the-art results for extracting MDAs from scientific publications [323]. The extracted MDAs will further expand the human MDA database. We expect that those deep learning methods will be more powerful with an expanded human MDA database.
@@ 3.4 Microbial Ecology
Microbial ecology seeks to decode the intricate dynamics and functional roles within complex microbial communities, from soil rhizospheres to the human gut. AI has emerged as a transformative force in this domain, enabling the extraction of predictive patterns from high-dimensional, noisy, and often sparse datasets that defy traditional statistical modeling. This section explores key applications where AI drives breakthroughs in microbial ecology: from predicting community composition to identifying keystone species, predicting community dynamics, community control, and source tracking. By integrating AI techniques with ecological theory, these applications not only enhance mechanistic understanding but also empower predictive stewardship of microbiomes in natural and engineered systems.
@@ 3.4.1 Microbial composition prediction
Being able to predict microbial community composition is a cornerstone of modern microbiome science. Microbial assemblages underpin essential ecological and physiological processes—from nutrient cycling in the environment to metabolism and immunity in host-associated systems—and their composition strongly determines overall community function. Predictive capability marks a shift from descriptive cataloging (“who is there") toward a mechanistic understanding of why and how communities assemble and change. Such models provide insight into the deterministic and stochastic forces shaping microbial ecosystems and enable the rational design and control of microbiomes for desired outcomes, including bioremediation, agricultural productivity, and human health interventions. In medicine, for example, accurate prediction of community compositions could inform diet-based therapies, probiotic design, and microbiota restoration after antibiotic treatment. Ultimately, the ability to forecast microbial community composition transforms microbiome research into a predictive, intervention-oriented discipline, akin to how meteorology evolved from observation to accurate weather forecasting.
Early attempts include a bioclimatic modeling approach that leverages neural networks to predict microbial community structure as a function of environmental parameters and microbial interactions [327], and the use of MLP to predict the temporal gut community composition of termite perturbed by six different lignocellulose food sources [328].
@@ 3.4.2 Keystone species identification
Keystone species are species whose influences on community structure and functioning are disproportionately large relative to its abundance. Microbial communities are also thought to contain keystone species, but applying macro-ecological approaches to large, complex microbiomes is challenging. Experimental removals are infeasible due to technical limitations and ethical concerns, particularly in host-associated systems such as the human gut microbiome. Statistical comparisons are similarly difficult, as it is rare to find two microbial communities differing by only a single species, and such comparisons are further complicated by strong inter-individual variation and numerous confounders. An alternative is to infer population-dynamics models and identify keystone species through simulated removals, but these methods require accurate model specification and high-quality absolute abundance data. Keystone species were also considered highly connected or high-betweenness taxa in microbial correlation networks. However, correlation-based edges capture statistical co-occurrence rather than direct ecological interactions, and network topology does not account for community specificity, e.g., a species may function as a keystone in one community but not in another.
@@
3.4.3 Microbial dynamics prediction
A fundamental question in microbial ecology is whether we can predict the temporal behaviors of complex microbial communities [332]. Traditionally, this problem is addressed using system identification or network reconstruction techniques, which assume specific population dynamics described by a set of ordinary differential equations. For example, the classical GLV model in community ecology, which considers pair-wise interactions, can be represented as a directed, signed, and weighted graph, often referred to as an ecological network. Numerous methods have been developed to infer these dynamics and reconstruct the ecological network using temporal or steady-state data [333]. However, this network-based approach typically assumes that inter-species interactions are exclusively pair-wise, which may not reflect the true nature of complex microbial interactions.
Recently, deep learning techniques have been deployed to predict temporal behaviors of microbiomes. For example, in 2022, Baranwal et al. applied LSTM (a classical variant of RNN) to learn from experimental data on temporal dynamics and functions of microbial communities to predict their future behavior and design new communities with desired functions [334]. Using a significant amount of experimental data, they found that this method outperforms the widely used GLV model in community ecology. In 2023, Thompson et al. proposed the Microbiome Recurrent Neural Network (MiRNN) architecture [167]. Inputs to the MiRNN at time step t1t-1 include the state of species abundances, metabolite concentrations, control inputs, and a latent vector that stores information from previous steps and whose dimension determines the flexibility of the model.The output from each MiRNN block is the predicted system state and the latent vector at the next time step $t$. To avoid the physically unrealistic emergence of previously absent species, a constrained feed-forward neural network outputs zero-valued species abundances if species abundances at the previous time step were zero. The authors demonstrated that MiRNN yielded comparable prediction performance to the LSTM model, but with more than a 50,000 fold reduction in the number of model parameters. These works are of broad interest to those working on microbiome prediction and design to optimize specific target functions. So far, LSTM and MiRNN have been just applied to synthetic communities with 25 diverse and prevalent human gut species and 4 major health-relevant metabolites (acetate, butyrate, lactate, and succinate). Its potential to large systems, e.g., the human gut microbiome, with thousands of species and metabolites would be interesting to explore. The quality of the training data would be crucial.
Autoencoders (AEs) have recently been employed to compress microbial growth curves into compact, low-dimensional embeddings [335]. These embeddings enable high-fidelity reconstruction of the original growth curves while preserving critical biological information despite substantial dimensionality reduction. Remarkably, they retain sufficient signal to distinguish complex phenotypes—such as strain identity and antibiotic resistance—predict microbial dynamics from initial conditions and experimental variables, and map directly to interpretable mechanistic growth parameters. In many cases, AE-derived embeddings outperform full growth curves on these tasks, indicating effective removal of extraneous noise without sacrificing biologically relevant information. Furthermore, AEs facilitate the compression of community dynamics into fewer variables than typically required by conventional mechanistic models.
In addition to methods specifically designed for predicting microbial dynamics, existing methodologies developed for multiple time series forecasting (MTSF) can also be potentially employed. For example, MTSF-DG is a model capable of learning historical relation graphs and predicting future relation graphs to capture dynamic correlations [168]. Evaluating the performance of these general time series prediction methods in the context of microbial dynamics prediction would be very interesting.